I have $69,000.00 to invest for one year. If B rate bonds pay 4 percent per year and CD's pay 13 percent per year, how much money in each fund do I have to invest to obtain $5730.00 after one year.
Answer by checkley71(8403)

5,730=.13x+.04(69,000-x) or 5,730=.13x+2,760-.04x or 5,730-2,760=.13x-.04x or 2,970=.09x or x=2,970/.09 or x=33,000 invested @ 13% & 69,000-33,000=36,000 invested @ 4%
proof 33,000*.13+36,000*.04=5,730 or 4,290+1,440=5,730 or 5,730=5,730